OpenStudy (anonymous):

Help please! Ivan walks to school one mile/hr faster than when he walks back home. If it takes him 12 min. (0.20 hr) to walk to school and 18 min. (0.30 hr) to walk back, how fast does he walk to school and how far does he have to walk to school?

OpenStudy (anonymous):

Set x is the speed for him to walk to school The distant from home - school: d d= 0.2x Speed from school -home: x-1 d= 0.3(x-1) 0.3(x-1)=0.2x because the same way 0.3x-0.3=0.2x 0.1x=0.3 X=3 mile/hr D=3.0.2=0.6mile

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

You are welcome:)